Ticket Number: HEA-NOC/20060817-1 Ticket Status: CLOSE
Ticket Type: unscheduled Resolver: ESB
Ticket Opened: 20060817 09:50 UTC+1 Problem Start: 20060817 08:20 UTC+1
Ticket Closed: 20060818 11:44 UTC+1 Problem End: 20060817 09:30 UTC+1
Loss of Connectivity to CIT
Traffic to and from CIT
20060817 09:44 OMG Monitoring tools indicated a problem reaching CIT this
morning. The problem seems to have resolved itself while I was
troubleshooting. I've contacted Diarmuid and Martin to see if they know of
any local issues that would have caused this.
20060817 09:51 OMG Martin responded saying that it appears that their
router was powered off.
20060817 10:54 OMG Diarmuid replied saying that there was an ESB outage in
the Bishopstown area which affected power to the campus. Will monitor for
24 hours before closing this ticket.
20060818 11:27 OMG No further problems have been observed. Closing ticket.
Time to Fix:
1 hour 10 minutes approx.
This ticket can be monitored at http://www.hea.net/tickets/20060817-1